An ideal gas with Cv = 1.62 kJ/kg-K and R = 273J/kg-K expands from 0.15 m3 and 27oC to 0.35 m3 during an isobaric process with p = 150kPa. Consider the process reversible nonflow, determine (a) the heat, (b) the change in internal energy, © the work nonflow, (d) the change in entropy. 


1
Expert's answer
2021-12-06T09:40:24-0500

a)

"T_1=300~K,"

"T_2=T_1\frac{p_2}{p_1}=700~K,"

"Q=\nu R\Delta T=53.2~kJ,"

b)

"\Delta U=p\Delta V=30~kJ,"

c)

"A=Q-\Delta U=23~kJ,"

d)

"S=\frac Q{\Delta T}=133~\frac JK."